# Search for Multiple Adjacent Marked Vertices on the Hypercube by a Quantum Walk with Partial Phase Inversion

### v2

@@ Abstract
There is a strong interest in quantum search algorithms, particularly in problems with multiple adjacent solutions. In the hypercube, part of the energy of the quantum system is retained in states adjacent to the target states, decreasing the chances of the target states being observed. This paper applies the Multiself-loop Lackadaisical Quantum Walk with Partial Phase Inversion to search for multiple adjacent marked vertices on the hypercube. Aspects like the type of marked vertices are considered in addition to using multiple self-loops and weight compositions. Two scenarios are analyzed. Firstly, the relative position of non-adjacent marked vertices together with adjacent marked vertices. Secondly, only adjacent marked vertices are analyzed. Here, we show experimentally that, with partial phase inversion, a quantum walk can amplify the probability amplitudes of the target states, reaching success probabilities of values close to $1$. We also show that the relative position of non-adjacent marked vertices does not significantly influence the search results. Our results demonstrate that the partial phase inversion of target states is a promising alternative to search adjacent solutions with quantum walks, which is a key capacity for real search applications.
Keywords Quantum Computing  $\cdot$ Quantum Walks  $\cdot$ Quantum Search Algorithm  $\cdot$ Lackadaisical Quantum Walk  $\cdot$ Multiple Self-loops  $\cdot$ Partial Phase Inversion  $\cdot$ Adjacent Marked Vertices
@@ 1 Introduction
Many advances have been achieved since the article’s publication by Aharonov et al. (1993), which is considered the first in quantum walks. One of the first quantum search algorithms based on quantum walks was designed by Shenvi et al. (2003), which defined the quantum walks as one of the most promising resources and an intuitive framework for building new quantum algorithms. Many other works on quantum walks have been developed since this moment (Ambainis et al., 2004; Potoček et al., 2009; Hein and Tanner, 2009; Ambainis et al., 2012).
@@
Amongst many proposed works in quantum walks, Wong (2015) developed a quantum search algorithm called lackadaisical quantum walks - LQW, an analog of classical lazy random walks in which the quantum walker has a chance to stay at the current vertex position introducing $m$ self-loops of integer weight $l$ at each vertex of the complete graph. This proposal was altered by Wong (2017) where the $m$ self-loops were reduced to one self-loop of non-integer weight. In turn, Souza et al. (2023) proposed a new quantum search algorithm based on the LQW called Multiself-Loop Lackadaisical Quantum Walk with Partial Phase Inversion - MSLQW – PPI, which uses $m$ self-lops in each vertex on the hypercube with weight value $l=l^{\prime}\cdot m$. The weight value $l^{\prime}\in\mathbb{R}$ and $m\in\mathbb{Z}$.
@@
However, some other studies indicate that the type of marked vertices influences the results of quantum search algorithms, particularly the adjacent marked vertices. According to Potoček et al. (2009), the final state of the algorithm designed by Shenvi et al. (2003) is mainly composed of the marked state but also part of the probability amplitude is retained in adjacent states. Another behavior of quantum walks on the hypercube referring to adjacent marked vertices is the formation of stationary states (Nahimovs et al., 2019). Souza et al. (2021) experimentally showed that adjacent marked vertices interfere with the search results. However, they have proposed a new ideal value of weight $l=(d/N)\cdot k$ when there are adjacent marked vertices in the set of solutions, a decrease in the maximum probability of success occurs.
@@
Therefore, this work objective is to apply MSLQW – PPI to research multiple marked vertices on the hypercube in two scenarios. The first scenario analyzes the research by multiple adjacent and no-adjacent vertices to verify that the relative position of non-adjacent vertices interferes with the search results. The second scenario examines the research by multiple adjacent vertices. The coefficient of variation was also used to evaluate the dispersion around the average maximum probability according to the relative position of the non-adjacent marked vertices. The results indicate insignificant variation around the maximum mean success probability. These results are significant because they show that the partial inversion of the target state based on multiple self-loops provides a new perspective of advances in developing new quantum search algorithms.
@@
This paper is organized as follows. Section 2 presents some concepts about Multiself-loop Lackadaisical Quantum Walks with Partial Phase Inversion on the hypercube. Section 3 the experiments are defined. Section 4 presents the results and discussion. Finally, in Section 5 are the conclusions.
2 MSLQW – PPI on the hypercube
@@
|$\ket{s^{C}}=\frac{1}{\sqrt{n+l}}\left(\sqrt{l^{\prime}}\sum_{j=0}^{m-1}\ket{\circlearrowleft_{j}}+\sum_{i=0}^{n-1}\ket{i}\right)$ (2)
and $l=l^{\prime}\cdot m$. The MSLQW – PPI system on the hypercube is started according to Equation 3.
|$\ket{\Psi(0)}=\frac{\sqrt{l^{\prime}}}{\sqrt{N}\times\sqrt{n+l}}\sum_{\vec{x}}\sum_{j=0}^{m-1}\ket{\vec{x},\circlearrowleft_{j}}+\frac{1}{\sqrt{N}\times\sqrt{n+l}}\sum_{\vec{x}}\sum_{i=0}^{n-1}\ket{\vec{x},i}$ (3)
Consider the quantum walk with search. A query to the “Grover oracle” is included in each step of the quantum walk $U^{\prime}=U\cdot(I_{n}\otimes Q)$ as follow,
$Q$ $=I_{(n+m)\cdot N}-2\sum_{\omega}\sum_{\epsilon=1}^{n}\ket{\epsilon,\omega}\bra{\epsilon,\omega}-2\sum_{\omega}\sum_{\tau}\ket{\circlearrowleft_{\tau},\omega}\bra{\circlearrowleft_{\tau},\omega}$ | (4)
where |$\ket{\omega}$ represents the marked vertex, $\epsilon$ represents an edge that is not a self-loop, and $\circlearrowleft_{\tau}$ are the self-loops that will have their phases inverted. The proposed modification of Grover’s oracle described in Equation 4 makes it possible to identify the components of the target state that have their phase inverted.
@@ 3 Experiment setup
The simulations performed in this work are divided into the following two scenarios. In the first scenario, we consider both adjacent and non-adjacent marked vertices. In the second scenario, we consider only the adjacent marked vertices. As in Souza et al. (2023), we flipped the phase of just a single self-loop, i.e., $\circlearrowleft_{\tau=0}$.
@@ 3.1 Definition of marked vertex samples
According to the definitions of the hypercube, two vertices are adjacent if the Hamming distance between them is $1$. Non-adjacent vertices have a Hamming distance of at least $2$ from any other vertex. We define the set of marked vertices divided into $M_{k,j}$ groups of samples with $k$ vertices and $j$ samples.
The adjacent and non-adjacent vertices form the first set. This set is divided into twelve groups of one hundred samples. For each sample of $k$ adjacent vertices, other $k-1$ non-adjacent vertices are marked, and thirty MSLQW – PPI are performed. Therefore, thirty-six hundred simulations are performed. For every hundred simulations, we preserve the same $k$ adjacent marked vertices and vary the locations of the $k-1$ non-adjacent marked vertices. For example, if $k=3$, we have two adjacent marked vertices and one non-adjacent marked vertex, $M_{3,100}=[\{0,1,1128\}_{1},\{0,1,2950\}_{2},\dots,\{0,1,1470\}_{100}]$.
The adjacent vertices form the second set. This set is divided into twelve sample groups containing between $2$ and $13$ marked vertices. To search for adjacent vertices, twelve simulations are performed. Initially, we have two marked vertices, and at each new simulation, a new vertex is marked and added to the new group as follows: $M_{2,1}=\{0,1\},\dots,M_{13,1}=\{0,1,2\dots,1024,2048\}$, until all adjacent vertices are marked.
The samples have $k$ distinct vertices, i.e., without replacement. The simulations performed in the first scenario set were necessary to obtain the average behavior based on the relative position of the non-adjacent marked vertices and verify their influence on the results. The stop condition for a simulation occurs after each of the thirty walks obtains the maximum value of the probability amplitude. In each quantum walk, a number $m$ of self-loops per vertex was defined, which varies between $1$ and $30$. The weight $l$ is distributed by dividing its value between $m$ equal parts.
@@ 3.2 Hardware and software setup for the simulations
The simulations were performed using the Parallel Experiment for Sequential Code - PESC Santos et al.  Henrique et al. (2023) to perform computational simulations distributed over a network. The programming language used to write the algorithms was Python 3.7. All machines used in the simulations utilize the operational system, Ubuntu 18.04.6 LTS (Bionic Beaver).
@@ 4 Results and discussion
As previously defined, the experiments are divided into two scenarios according to the type of marked vertices. In the first scenario, we have adjacent and non-adjacent marked vertices. As we analyzed the relative positional of the non-adjacent marked vertices, thirty-six thousand simulations were performed, which were divided into twelve groups with one hundred samples from k vertices, and to each sample was performed thirty quantum walks MSLQW – PPI. Then, the variability of the results was also analyzed and is represented in Fig. 2. In the second scenario, we only have adjacent marked vertices. Each node has the same number of adjacent vertices as the hypercube’s degree number. Therefore, twelve simulations were realized, and thirty MSLQW – PPI were performed in each simulation. The results are represented in Figures 1 and 3 respectively. They present the maximum probability of success according to the number of self-loops and marked vertices.
@@ 4.1 Analyzing the search with adjacent and non-adjacent marked vertices
Fig. 1a shows the probability of success for the weight $l=n/N$. Rhodes and Wong (2020) proposed this weight value to search a single vertex, while Souza et al. (2021) used it to search multiple vertices. However, the results showed that this weight value is not ideal. Souza et al. (2023) used this weight value and applied MSLQW – PPI to search for multiple non-adjacent marked vertices, but there was no increase in the maximum probability of success. In this article, the maximum average probability obtained $p=0.999$ with three marked vertices (two adjacent vertices and one non-adjacent vertex) and a single self-loop, which is a result close to that achieved by Souza et al. (2021) of $p=0.997$. In both cases, the maximum probability of success decreases as the number of marked vertices increases.
@@
Fig. 1b shows the success probability using the weight $l=(n/N)\cdot k$. In cases with only non-adjacent marked vertices, for this weight value, only a single self-loop is needed (Souza et al., 2021, 2023). Although, when there are adjacent marked vertices, it is necessary to increase the number of self-loops and use partial phase inversion to obtain success probabilities close to $1$, as we can see in Table 1. In some cases, we can observe an improvement in the probability of success compared to the results obtained using only one self-loop.
Figure 1: The probability of success of the MSLQW – PPI to search for adjacent and non-adjacent marked vertices with $n=12$ and $N=4096$ vertices. LABEL:sub@fig:probability-distribution-neighbors-non-neighbors-a weight value $l=n/N$. LABEL:sub@fig:probability-distribution-neighbors-non-neighbors-b weight value $l=(n/N)\cdot k$. LABEL:sub@fig:probability-distribution-neighbors-non-neighbors-c weight value $l=n^{2}/N$. LABEL:sub@fig:probability-distribution-neighbors-non-neighbors-d weight value $l=(n^{2}/N)\cdot k$.
@@
Comparing columns A and C of Table 1, we can see that using partial phase inversion from a certain quantity of self-loops, it is possible to obtain more significant probabilities than those achieved with the use of a single self-loop. Now, comparing columns B and C with at least two self-loops, it is possible to improve the maximum probability of success. However, for this self-loop weight value, as the number of marked vertices increased, only a single self-loop is needed to achieve probabilities of approximately $p\approx 0.98$.
Fig. 1c shows the probability of success using the weight value $l=n^{2}/N$. This weight is proposed by Souza et al. (2023) and is composed of the weight value presented by Rhodes and Wong (2020) to search for one marked vertex plus an exponent in the element that represents the degree of the vertex in the numerator. Compared with the results found by Souza et al. (2021) and with the results shown in Fig. 1a, there was a significant improvement in the maximum probability of success for numbers $k>3$ marked vertices. In this scenario, the success probabilities depend on the inversely proportional relationship between the number $k$ of marked vertices and the number $m$ of Self-loops. This means that as the number of marked vertices increases, the number of self-loops decreases and the other way around, however, the maximum probability of success continues above $p=0.97$.
@@
Table 1: Cases for searching adjacent and non-adjacent marked vertices where more than one self-loop is required to obtain a maximum probability close to $1$ using the weight $l=(n/N)\cdot k$ proposed by Souza et al. (2021).

### v3

@@ Abstract
Quantum walks provide an efficient tool for the construction of new quantum search algorithms. In this paper, we analyze the application of the Multiself-loop Lackadaisical Quantum Walk on the hypercube that uses partial phase inversion of the target state to search for multiple adjacent marked vertices. We consider two scenarios and one of them evaluates the influence of the relative position of non-adjacent marked vertices on the search results. The use of self-loops and the composition of their weights are an essential part of the construction process of new quantum search algorithms based on lackadaisical quantum walks, however, other aspects have been considered, such as, for example, the type of marked vertices. It is known that part of the energy of a quantum system is retained in states adjacent to the target state. This behavior causes the amplification of these states where the sum of probability amplitudes reaches values equivalent to those of the target state, reducing their chances of being observed. Here we show experimentally that with the use of partial phase inversion of the target state, it is possible to amplify its probability amplitudes even in scenarios with adjacent marked vertices reaching maximum success probabilities with values close to $1$. We also show that the relative position of the non-adjacent marked vertices did not significantly influence the results. The lackadaisical quantum walk generalization in using multiple self-loops to only a single self-loop and the ideal composition of a weight value was sufficient to obtain advances to quantum search algorithms based on quantum walks. However, the results presented here show that many other aspects need to be taken into account for the construction of new quantum algorithms. It was possible to add gains in the maximum probabilities of success compared to other results found in the literature. In one of the most significant cases, the probability of success increased from $p\approx 0.38$ to $p>0.99$. Therefore, the use of partial phase inversion of target states brings new contributions to the development of new quantum search algorithms based on quantum walks and the use of multiple self-loops.
Keywords Quantum Computing  $\cdot$ Quantum Walks  $\cdot$ Quantum Search Algorithm  $\cdot$ Lackadaisical Quantum Walk  $\cdot$ Multiple Self-loops  $\cdot$ Adjacent Marked Vertices  $\cdot$ Partial Phase Inversion
@@ 1 Introduction
Many advances have been achieved since the publication of the article by Aharonov et al. (1993), which is considered the first in quantum walks. One of the first quantum search algorithms based on quantum random walks was designed by Shenvi et al. (2003), which defined the quantum walks as one of the most promising resources and an intuitive framework for building new quantum algorithms. Many other works on quantum walks have been developed since this moment (Ambainis et al., 2004; Potoček et al., 2009; Hein and Tanner, 2009; Ambainis et al., 2012).
@@
Amongst many proposed works in quantum walks, Wong (2015) developed a quantum search algorithm called lackadaisical quantum walks - LQW, an analog of classical lazy random walks in which the quantum walker has a chance to stay at the current vertex position by introducing $m$ self-loops of integer weight $l$ at each vertex of the complete graph. This proposal was altered by Wong (2017) where the $m$ self-loops were generalized to one self-loop of non-integer weight. In turn, Souza et al. (2023) proposed a new quantum search algorithm based on the LQW called Multiself-Loop Lackadaisical Quantum Walk - MSLQW, which uses $m$ self-lops in each vertex on the hypercube with weight value $l=l^{\prime}\cdot m$, and the partial phase inversion of the target state to research multiple marked vertices. The weight value $l^{\prime}\in\mathbb{R}$ and $m\in\mathbb{Z}$.
@@
However, some other studies indicate that the type of marked vertices influences the results of quantum search algorithms, in particular, the adjacent marked vertices. According to Potoček et al. (2009), the final state of the algorithm designed by Shenvi et al. (2003) is mainly composed of the marked state but also has small contributions from its nearest neighbors, i.e., part of the probability amplitude is retained in adjacent vertices. Another behavior of quantum walks on the hypercube referring to adjacent marked vertices is the formation of stationary states (Nahimovs et al., 2019). Souza et al. (2021) experimentally showed that adjacent marked vertices interfere with the results of the search for multiple marked vertices. Although they have proposed a new ideal value of weight $l=(d/N)\cdot k$, when there are adjacent marked vertices in the set of solutions occurs a decrease in the maximum probability of success.
Therefore, this work objective is to apply MSLQW-PPI to research multiple marked vertices in two scenarios. The first scenario analyzes research by multiple adjacent and no-adjacent vertices to verify that the relative position of non-adjacent vertices interferes with the search results. The second scenario analyzes research by multiple adjacent vertices. Based on the methodology used by Souza et al. (2023), the results presented in this work are promising. Comparing the results of this work with the results obtained by Souza et al. (2021) there was a gain in the maximum probability of success to values close to $1$. Before some of the success probabilities reached only $p\approx 0.59$ and $p\approx 0.38$, the first and second scenarios respectively.
@@
In their proposal, Souza et al. (2023) used the search for non-adjacent marked vertices with the phases of $1\leqslant s<m$ inverted self-loops and $1\leqslant m\leqslant 30$. The results indicated that, by inverting the phase of only one self-loop, it is possible to achieve maximum probabilities of success close to $1$. Based on the results of Souza et al. (2023) we applied the particular case of MSLQW-PPI with $s=1$ inverted self-loop. Compared to the results obtained in this work, the maximum success probabilities remain close to $1$. The coefficient of variation was also used to evaluate the dispersion around the average relative position of the non-adjacent marked vertices. The coefficient of variation was also used to evaluate the dispersion around the average maximum probability according to the relative position of the non-adjacent marked vertices. The results indicate that the variation around the maximum mean probability of success is not significant. These results are important because they show that the partial inversion of the target state based on the use of multiple self-loops provides a new perspective of advances in the development of new quantum search algorithms.
@@
This paper is organized as follows. In Section 2 we present some concepts about Multiself-loop Lackadaisical Quantum Walks in the hypercube. Section 3 the experiments are defined. Section 4 presents the results and discussion. Finally, in Section 5 are the conclusions.
2 Multi-self-loop lackadaisical quantum walk on the hypercube
The lackadaisical quantum walk is the quantum counterpart of the classical lazy random walk. This quantum algorithm was proposed by Wong (2015) and is obtained by adding a self-loop to each vertex of the graph. Then, the Hilbert space associated with the lackadaisical quantum walk in the hypercube is
$\mathcal{H}=\mathcal{H}^{n+1}\otimes\mathcal{H}^{2^{n}}$
where $\mathcal{H}^{n+1}$ is the Hilbert space associated with the quantum coin space, and $\mathcal{H}^{2^{n}}$ is the Hilbert space associated with nodes in the hypercube. According to Høyer and Yu (2020), in a $n$-regular graph by adding a self-loop of weight $l$ to each vertex, the coined Hilbert space becomes
$\mathcal{H}^{n+1}=\{\ket{e_{0}},\ket{e_{1}},\dots,\ket{e_{n-1}},\ket{\circlearrowleft}\}.$
where $e_{i}$ is a binary string of $n$ bits with $1$ in the $i$-th position (Kempe, 2002; Shenvi et al., 2003), and $\ket{\circlearrowleft}$ is the self-loop. Weighted self-loop accounting is done by modifying Grover’s coin as follows
$C=2\ket{s^{C}}\bra{s^{C}}-I_{(n+1)}$ (1)
where
|$\ket{s^{C}}=\frac{1}{\sqrt{n+l}}\left(\sqrt{l}\ket{\circlearrowleft}+\sum_{i=0}^{n-1}\ket{i}\right).$ (2)
The Lackadaisical Quantum Walk system in the hypercube starts as follows
|$\ket{\Psi(0)}=\frac{1}{\sqrt{N}}\sum_{\vec{x}}\ket{\vec{x}}\otimes\ket{s^{C}}.$ (3)
Substituting Equation 2 into Equation 3 we obtain the initial state described in Equation 4.
|$\ket{\Psi(0)}=\frac{\sqrt{l}}{\sqrt{N}\times\sqrt{n+l}}\sum_{\vec{x}}\ket{\vec{x},\circlearrowleft}+\frac{1}{\sqrt{N}\times\sqrt{n+l}}\sum_{\vec{x}}\sum_{i=0}^{n-1}\ket{\vec{x},i}$ (4)
The Multiself-loop lackadaisical quantum walk was proposed by Souza et al. (2023). This quantum algorithm is obtained by adding $m$ self-loops at each vertex of the hypercube and a partial phase inversion of the target state is applied. The Hilbert space associated with the lackadaisical quantum walk in the hypercube is
$\mathcal{H}=\mathcal{H}^{n+m}\otimes\mathcal{H}^{2^{n}}.$
Then, the Hilbert space associated with the coin space becomes
@@ 2 Multi-self-loop lackadaisical quantum walk on the hypercube
To account for the weighted auto-loop, a modification is made to the Grover coin described in Equation 1 as follows
@@
|$\ket{s^{C}}=\frac{1}{\sqrt{n+l}}\left(\sqrt{l^{\prime}}\sum_{j=0}^{m-1}\ket{\circlearrowleft_{i}}+\sum_{i=0}^{n-1}\ket{i}\right)$ (6)
and $l=l^{\prime}\cdot m$. The Multiself-loop lackadaisical quantum walk system on the hypercube is also started according to Equation 3. Substituting Equation 6 into Equation 3, we obtain the initial state described in Equation 7.
|$\ket{\Psi(t=0)}=\frac{\sqrt{l^{\prime}}}{\sqrt{n+l}\times\sqrt{N}}\sum_{j=0}^{m-1}\sum_{\vec{x}}\ket{\circlearrowleft_{j},\vec{x}}+\frac{1}{\sqrt{n+l}\times\sqrt{N}}\sum_{i=0}^{n-1}\sum_{\vec{x}}\ket{i,\vec{x}}$ (7)
The proposed modification of the oracle described in Equation 8 makes it possible to identify the components of the target state.
$Q=I_{(n+\text{m})\cdot N}-2\sum_{\omega}\sum_{\epsilon=1}^{n}\ket{\epsilon,\omega}\bra{\epsilon,\omega}-2\sum_{\omega}\sum_{\tau}\ket{\circlearrowleft_{\tau},\omega}\bra{\circlearrowleft_{\tau},\omega}$ | (8)
where |$\ket{\omega}$ represents the marked vertex, $\epsilon$ represents an edge that is not a self-loop, and $\circlearrowleft_{\tau}$ is the self-loop that will have its phase inverted. Consider an arbitrary state which denotes the superposition of all edges
[…]
Consider the states |$\ket{\circlearrowleft_{\tau}}$ as the target self-loops and $s=1$. Applying the phase inversion operator, represented by Equation 10.
$Q=I_{(n+\text{m})\cdot N}-2\sum_{\omega}\sum_{\epsilon=0}^{n-1}\ket{\epsilon,\omega}\bra{\epsilon,\omega}\\ -2\sum_{\omega}\sum_{\tau}\ket{\circlearrowleft_{\tau=0},\omega}\bra{\circlearrowleft_{\tau=0},\omega}$ | (10)
where $I_{n+m}$ is the identity operator of dimension $n+m$ and |$\ket{\circlearrowleft_{\tau=0}}$ as the target self-loop, we have
[…]
@@ 3 Experiment setup
According to Shenvi et al. (2003); Potoček et al. (2009), part of the probability amplitude of a quantum walk on the hypercube is retained at vertices adjacent to a marked vertex, and if two marked vertices on the hypercube are adjacent stationary states are formed (Nahimovs, 2019). Souza et al. (2021) showed that adjacent marked vertices interfere with the search performance. Therefore, the experiments performed in this work are divided into the following two scenarios. In the first scenario, we consider both adjacent and non-adjacent marked vertices. In the second scenario, we consider only the adjacent marked vertices.
@@ 3.1 Definition of marked vertex samples
According to the definitions of the hypercube, two vertices are adjacent if the Hamming distance between them is $1$. Non-adjacent vertices are those that have a Hamming distance of at least $2$ from any other vertex. In this way, we define the set of marked vertices to execute the simulations. The set of marked vertices is divided into $M_{k,\gamma}$ groups of samples with $k$ vertices and $j$ samples.
The first set is formed by the vertices that are both adjacent and non-adjacent. This set is divided into twelve groups of one hundred samples. For each sample of $k$ adjacent vertices, other $k-1$ non-adjacent vertices are also marked, and thirty MSLQW-PPI are performed. Therefore, thirty-six hundred simulations are performed. Every hundred simulations we preserve the same $k$ adjacent marked vertices and vary the locations of the $k-1$ non-adjacent marked vertices, for example, if $k=3$ we have two adjacent marked vertices and one non-adjacent marked vertex,
$M_{3,100}=[\{0,1,1128\}_{1},\{0,1,2950\}_{2},\dots,\{0,1,1470\}_{100}]$
Every one hundred new simulations, $k$ new adjacent vertices and $(k-1)\cdot 100$ new non-adjacent vertices are marked and added to the group until $k+(k-1)=5,7,\dots,25$.
The second set is formed by the adjacent vertices. This set is divided into twelve groups of one sample that contain between $2$ and $13$ marked vertices. To search for adjacent vertices, twelve simulations are performed. Initially, we have two marked vertices and at each new simulation a new vertex is marked and added to the new group as follows
$M_{2,1}=\{0,1\},M_{3,1}=\{0,1,2\},\dots,M_{13,1}=\{0,1,2,\dots,1024,2048\}$
until all adjacent vertices are marked.
The samples have $k$ distinct vertices, i.e., without replacement. The simulations performed in the set of the first scenario were necessary so that we could obtain the average behavior based on the relative position of the non-adjacent marked vertices and verify their influence on the results. In each simulation, thirty MSLQW-PPI are performed. The stop condition for a simulation occurs after each of the thirty walks obtains the maximum value of the probability amplitude. In each quantum walk, a number $m$ of self-loops per vertex was defined, which varies between $1$ and $30$. The weight $l$ is distributed by dividing its value between $m$ equal parts.
@@ 3.2 Hardware and software setup for the simulations
The simulations were performed using the Parallel Experiment for Sequential Code - PESC (Henrique et al., 2023), to perform computational simulations distributed over a network. The platform provides a web interface for configuring simulation requests and manages the status and lifecycle of the request. The use of the platform simplified the simulations execution process, which was important to support the collect the study data. The tool is being developed for the instrumentation and optimization of the research group’s computational experiments. The programming language used to write the algorithms was Python 3.7. All machines that were used in the simulations utilize the operational system, Ubuntu 18.04.6 LTS (Bionic Beaver), and have an HD of 500 GiB. Table 2 shows the machines’ settings.
Table 1: Machine hardware configuration.
Table 2: *
Adapted from: (Souza et al., 2023). Machines System RAM System Processor 1 8 GiB Intel(R) Core(TM) i7-2600 CPU @ 3.40GHz 1 16 GiB Intel(R) Core(TM) i7-8700 CPU @ 3.20GHz 2 32 GiB Intel(R) Core(TM) i7-2600K CPU @ 3.40GHz 2 32 GiB Intel(R) Core(TM) i7-6700 CPU @ 3.40GHz
@@ 4 Results and discussion
As previously defined, the experiments are divided into two scenarios according to the type of marked vertices. In the first scenario, we have adjacent and non-adjacent marked vertices. As we analyzed the relative positional of the non-adjacent marked vertices, thirty-six thousand simulations were performed, which are divided into twelve groups with one hundred samples from $k$ vertices, and to each sample was performed thirty quantum walks MSLQW. Then, the variability of the results was also analyzed and is represented in Fig. 2. In the second scenario, we only have adjacent marked vertices. Each node has the same number of adjacent vertices as the hypercube’s degree number, therefore, twelve simulations were realized, and in each simulation, thirty MSLQW-PPI were also performed. The quantum walks vary according to the number of self-loops from one to thirty. The results are represented in Figures 1 and 3 respectively. They present the maximum probability of success according to the number of self-loops and marked vertices.
@@ 4.1 Analyzing the search with adjacent and non-adjacent marked vertices
Fig. 1a shows the probability of success for the weight $l=n/N$. Rhodes and Wong (2020) proposed this weight value to search a single vertex while Souza et al. (2021) used it to search multiple vertices, however, the results showed that this weight value is not ideal in this case. Souza et al. (2023) used this weight value and applied MSLQW-PPI to search for multiple non-adjacent marked vertices but there was no increase in the maximum probability of success. In this article, the maximum average probability obtained $p=0.999$ with three marked vertices (two adjacent vertices and one non-adjacent vertex) and a single self-loop, which is a result close to that achieved by Souza et al. (2021) of $p=0.997$. In both cases, as the number of marked vertices increases, the maximum probability of success decreases.
@@
Fig. 1b shows the success probability using the weight $l=(n/N)\cdot k$. In cases where there are only non-adjacent marked vertices, for this weight value, only a single self-loop is needed (Souza et al., 2021, 2023). Although, when there are adjacent marked vertices, it is necessary to increase the number of self-loops to obtain success probabilities close to $1$ as we can see in Table 3. In some cases, we can observe that there was an improvement in the probability of success compared to the results obtained with the use of only one self-loop.
@@
Comparing columns A and C of Table 3, we can see that from a certain quantity of self-loops, it is possible to obtain more significant probabilities than those achieved with the use of a single self-loop. It is necessary to use the least number of self-loops to obtain these results. Now, comparing columns B and C with at least two self-loops, it is possible to improve the maximum probability of success. However, for this self-loop weight value, as the number of marked vertices increased, only a single self-loop is needed to achieve probabilities of approximately $p\approx 0.98$.
(a) $l=n/N$
(b) $l=(n/N)\cdot k$
(c) $l=n^{2}/N$
(d) $l=(n^{2}/N)\cdot k$
Figure 1: The probability of success of the MSLQW-PPI to search for adjacent and non-adjacent marked vertices with $n=12$ and $N=4096$ vertices. LABEL:sub@fig:probability-distribution-neighbors-non-neighbors-a weight value $l=n/N$. LABEL:sub@fig:probability-distribution-neighbors-non-neighbors-b weight value $l=(n/N)\cdot k$. LABEL:sub@fig:probability-distribution-neighbors-non-neighbors-c weight value $l=n^{2}/N$. LABEL:sub@fig:probability-distribution-neighbors-non-neighbors-d weight value $l=(n^{2}/N)\cdot k$.
@@
Table 3: Cases for searching adjacent and non-adjacent marked vertices where more than one self-loop is required to obtain a maximum probability close to $1$ using the weight $l=(n/N)\cdot k$ proposed by Souza et al. (2021). Rows with ( - ) in column A mean that the same values are reached in the same rows in column B.